OpenStudy (anonymous):

I really need help; Let F(x)=8x-7 and g(x)=8x+8. Find (f*g)(7)

OpenStudy (anonymous):

any idea

OpenStudy (imstuck):

Find the f function's value at x=7 and then find the g function's value at x = 7. Then multiply those results together to get the answer.

OpenStudy (imstuck):

Any idea?

OpenStudy (anonymous):

i got lost can you explain it step by step

OpenStudy (anonymous):

i know the answer is 505 but i want to know the steps so i can really understand it

OpenStudy (aum):

(f*g)(7) = f(7) * g(7) fx) = 8x - 7 f(7) = ? g(x)=8x+8 g(7) = ? f(7) * g(7) = ?

OpenStudy (aum):

Is it (f*g)(7) OR (fog)(7) ?

OpenStudy (anonymous):

it is (f times g)

OpenStudy (aum):

No it is not!

OpenStudy (anonymous):

(f times g) (7)

OpenStudy (aum):

\[ (f \circ g)(7) \]

OpenStudy (anonymous):

yes

OpenStudy (aum):

That is not f times g. \[(f \circ g)(7) = f(g(7)) \\ g(x) = 8x + 8 \\ g(7) = ? \]

OpenStudy (anonymous):

f(x)=8x-7 and g(x)8x+8 Find (f∘g) (7)

OpenStudy (anonymous):

i still dont understand how to do the step

OpenStudy (aum):

Yes, you have to do it in two steps: (fog)(7) = f(g(7)) Find g(7) first. What do you get?

OpenStudy (aum):

g(x) = 8x + 8 g(7) = ? Can you answer this step first?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 64

OpenStudy (aum):

correct. \[ f(x) = 8x - 7 \\ f(g(7)) = f(64) = ? \]

OpenStudy (aum):

You calculated g(7) = 64 So f(g(7)) = f(64) = ?

OpenStudy (anonymous):

this is the part when i get lost

OpenStudy (aum):

f(x) = 8x - 7 f(64) = 8*64 - 7 = 512 - 7 = 505

OpenStudy (aum):